# A note on subgroups whose quotients are quasi-lines

### v1

@@ 4 Reducing a Euclidean cube complex to a line
If, in addition to the conditions of the proposition, one assumes for example that $\Gamma$ acts on $Y$ without a fixed point in the visual boundary then [Caprace-Sageev, Theorem E] implies that $\Gamma$ fixes a Euclidean flat in $Y$. The task of finding an appropriate map to $\mathbb{Z}$ then requires some work (and probably some CAT(0) geometry). It is moreover possible to reduce to the case where either there is no fixed point in the visual boundary or there is a fixed point in the Roller boundary [CFI, Proposition 2.26]. In the later case, Caprace’s appendix in [CFI] provides several epimorphisms from a finite index subgroup of $\Gamma$ to $\mathbb{Z}$ (in fact, in either case one can show that $\Gamma$ virtually fixes the simplicial boundary of $Y$ pointwise which suffices to apply Caprace’s argument). The fact that one of these has a kernel with the required properties is not immediate, but it can be shown following much the same argument as the proof of Proposition 1.7.
@@ 5 A family of examples
A family of groups introduced by Houghton [Houghton78] provides an example, for each $n\geq 2$, of a finitely generated pair of groups $H_{n}\leq G_{n}$ such that $H_{n}\operatorname{\setminus}G_{n}$ is an $n$-ended quasi-tree but $H_{n}$ is not a virtual fiber subgroup. These groups are known to provide examples of pairs of groups with an arbitrary number of relative ends [Scott77, Example 2.1] and, as will become apparent, the fact that $H_{n}\operatorname{\setminus}G_{n}$ is not only $n$-ended but a quasi-tree is obvious from the construction.
@@
Let $n\in\mathbb{N}$ and $X_{n}$ be a disjoint union of $n$ copies of $\mathbb{N}$. We will use the notation $X_{n}=\{1,\dots,n\}\times\mathbb{N}$. A bijection $f:X_{n}\rightarrow X_{n}$ is a translation at infinity of $X_{n}$ if there exists a finite subset $K\subseteq X_{n}$ and constants $a_{1},\dots,a_{n}\in\mathbb{Z}$ such that $f(i,m)=(i,m+a_{i})$ for all $(i,m)\in X_{n}-K$.
Houghton’s group $\mathcal{G}_{n}$ is the group of all translations at infinity of $X_{n}$.
When $n=1$, $\mathcal{G}_{n}$ is the group of finitely supported permutations on $\mathbb{N}$, which is not finitely generated. However $\mathcal{G}_{n}$ is finitely generated for all $n\geq 2$. For $n\geq 3$, Houghton defined these groups via the generating sets $S_{n}$ defined below. The equivalence of the two definitions essentially follows from the observation of Wiegold [Wiegold] that the commutator subgroup $[\operatorname{\langle}S_{n}\operatorname{\rangle},\operatorname{\langle}S_{n}\operatorname{\rangle}]$ is the group of finitely supported permutations of $X_{n}$ (see also [Lee, Lemma 2.7]) and is implicit in [Brown]. The generating sets are defined as follows.
Let $X$ be the increasing union $\cup_{n\in\mathbb{N}}X_{n}=\mathbb{N}\times\mathbb{N}$ and note that each $\mathcal{G}_{n}$ is a subgroup of the permutation group of $X$. For each $n\in\mathbb{N}$, let $g_{i}:X\rightarrow X$ be defined by
@@
Then $\mathcal{G}_{2}$ is generated by $S_{2}\coloneqq\{g_{1},\beta\}$ (see [Lee, pages 5-6]).
@@
Brown showed in [Brown, Theorem 5.1] that each $\mathcal{G}_{n}$ is of type $FP_{n-1}$ but not of type $FP_{n}$ for all $n$ and that $\mathcal{G}_{n}$ is finitely presented for all $n\geq 3$. Recall that, for any $n$, if a group $G$ is of type $FP_{n}$ and is finitely presented then $G$ is of type $F_{n}$ (see e.g. [Brown94, Section 8.7]). This allows us to construct, for any $n\geq 2$, a pair of groups $H\leq G$ such that both $G$ and $H$ are of type $F_{n}$ and the quotient space $H\operatorname{\setminus}G$ is a quasi-line but $H$ is not a virtual fiber subgroup.
@@
Fix $n\geq 3$ and let $\sigma$ be a permutation of $\{1,\dots,n\}$. Define $\alpha:X_{n}\rightarrow X_{n}$ by $\alpha(i,m)=(\sigma(i),m)$ and let $L$ be the subgroup of the permutation group of $X_{n}$ generated by $S_{n}\cup\{\alpha\}$. Note that $\alpha\mathcal{G}_{n}\alpha^{-1}=\mathcal{G}_{n}$ and $\alpha\notin\mathcal{G}_{n}$ unless $\sigma=\operatorname{id}$. Therefore $L=\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}$, where $m$ is the order of $\sigma$, and $L$ is of type $F_{n-1}$ by [Alonso, Corollary 9]. Let $M\coloneqq\operatorname{Stab}_{L}((1,1))$. Then $\alpha\in M$ so $M=H_{n}\rtimes\operatorname{\langle}\alpha\operatorname{\rangle}\cong\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}=L$. Moreover, the Schreier graph of $M$ with respect to $S_{n}\cup\{\alpha\}$ is a quasi-tree whose number of ends is the number of orbits of the action of $\operatorname{\langle}\sigma\operatorname{\rangle}$ on $\{1,\dots,n\}$ and $L=M\cup Mg_{1}M$.

### v2

@@ 5 Reducing a Euclidean cube complex to a line
If, in addition to the conditions of the proposition, one assumes for example that $\Gamma$ acts on $Y$ without a fixed point in the visual boundary then [CS11, Theorem E] implies that $\Gamma$ fixes a Euclidean flat in $Y$. The task of finding an appropriate map to $\mathbb{Z}$ then requires some work (and probably some CAT(0) geometry). It is moreover possible to reduce to the case where either there is no fixed point in the visual boundary or there is a fixed point in the Roller boundary [CFI16, Proposition 2.26]. In the later case, Caprace’s appendix in [CFI16] provides several epimorphisms from a finite index subgroup of $\Gamma$ to $\mathbb{Z}$ (in fact, in either case one can show that $\Gamma$ virtually fixes the simplicial boundary of $Y$ pointwise which suffices to apply Caprace’s argument). The fact that one of these has a kernel with the required properties is not immediate, but it can be shown following much the same argument as the proof of Proposition 1.11.
@@ 6 A family of examples
A family of groups introduced by Houghton [Hou78] provides an example, for each $n\geq 2$, of a finitely generated pair of groups $H_{n}\leq G_{n}$ such that $H_{n}\operatorname{\setminus}G_{n}$ is an $n$-ended quasi-tree with linear growth – and therefore narrow by Proposition 3.2 – but $H_{n}$ is not a virtual fiber subgroup. These groups are known to provide examples of pairs of groups with an arbitrary number of relative ends [Sco77, Example 2.1] and, as will become apparent, the fact that $H_{n}\operatorname{\setminus}G_{n}$ is not only $n$-ended but a quasi-tree with linear growth is obvious from the construction.
@@
Let $n\in\mathbb{N}$ and $X_{n}$ be a disjoint union of $n$ copies of $\mathbb{N}$. We will use the notation $X_{n}=\{1,\dots,n\}\times\mathbb{N}$. A bijection $f:X_{n}\rightarrow X_{n}$ is a translation at infinity of $X_{n}$ if there exists a finite subset $K\subseteq X_{n}$ and constants $a_{1},\dots,a_{n}\in\mathbb{Z}$ such that $f(i,m)=(i,m+a_{i})$ for all $(i,m)\in X_{n}-K$. Houghton’s group $\mathcal{G}_{n}$ is the group of all translations at infinity of $X_{n}$.
When $n=1$, $\mathcal{G}_{n}$ is the group of finitely supported permutations on $\mathbb{N}$, which is not finitely generated. However $\mathcal{G}_{n}$ is finitely generated for all $n\geq 2$. For $n\geq 3$, Houghton defined these groups via the generating sets $S_{n}$ defined below. The equivalence of the two definitions essentially follows from the observation of Wiegold [Wie77] that the commutator subgroup $[\operatorname{\langle}S_{n}\operatorname{\rangle},\operatorname{\langle}S_{n}\operatorname{\rangle}]$ is the group of finitely supported permutations of $X_{n}$ (see also [Lee12, Lemma 2.7]) and is implicit in [Bro87]. The generating sets are defined as follows.
Let $X$ be the increasing union $\cup_{n\in\mathbb{N}}X_{n}=\mathbb{N}\times\mathbb{N}$ and note that each $\mathcal{G}_{n}$ is a subgroup of the permutation group of $X$. For each $i\in\mathbb{N}$, let $g_{i}:X\rightarrow X$ be defined by
@@
Then $\mathcal{G}_{2}$ is generated by $S_{2}\coloneqq\{g_{1},\beta\}$ (see [Lee12, pages 5-6]).
@@
Brown showed in [Bro87, Theorem 5.1] that each $\mathcal{G}_{n}$ is of type $FP_{n-1}$ but not of type $FP_{n}$ for all $n$ and that $\mathcal{G}_{n}$ is finitely presented for all $n\geq 3$. Recall that, for any $n$, if a group $G$ is of type $FP_{n}$ and is finitely presented then $G$ is of type $F_{n}$ (see e.g. [Bro94, Section 8.7]). This allows us to construct, for any $n\geq 2$, a pair of groups $H\leq G$ such that both $G$ and $H$ are of type $F_{n}$ and the quotient space $H\operatorname{\setminus}G$ is a quasi-line but $H$ is not a virtual fiber subgroup.
@@
Fix $n\geq 3$ and let $\sigma$ be a permutation of $\{1,\dots,n\}$. Define $\alpha:X_{n}\rightarrow X_{n}$ by $\alpha(i,m)=(\sigma(i),m)$ and let $L$ be the subgroup of the permutation group of $X_{n}$ generated by $S_{n}\cup\{\alpha\}$. Note that $\alpha\mathcal{G}_{n}\alpha^{-1}=\mathcal{G}_{n}$ and $\alpha\notin\mathcal{G}_{n}$ unless $\sigma=\operatorname{id}$. Therefore $L=\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}$, where $m$ is the order of $\sigma$, and $L$ is of type $F_{n-1}$ by [Alo94, Corollary 9]. Let $M\coloneqq\operatorname{Stab}_{L}((1,1))$. Then $\alpha\in M$ so $M=H_{n}\rtimes\operatorname{\langle}\alpha\operatorname{\rangle}\cong\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}=L$. Moreover, the Schreier graph of $M$ with respect to $S_{n}\cup\{\alpha\}$ is a quasi-tree whose number of ends is the number of orbits of the action of $\operatorname{\langle}\sigma\operatorname{\rangle}$ on $\{1,\dots,n\}$ and $L=M\cup Mg_{1}M$.
@@
[…]
Fix a (set-theoretic) section $\varphi:Q\rightarrow G$ of $\pi$. If $K\leq Q$ is a finitely generated subgroup such that $K\operatorname{\setminus}Q$ is an $n$-ended quasi-tree for some $2\leq n<\infty$ then the pull-back $H\coloneqq\pi^{-1}(K)$ is a finitely generated subgroup of $G$ and $H\operatorname{\setminus}G$ is an $n$-ended quasi-tree. Moreover, if there is a finite subset $F\subseteq Q$ such that $Q=KFK$ then $G=H\varphi(F)H$. Therefore if $K\leq Q$ is not a virtual fiber then neither is $H\leq G$. As demonstrated by Example 6.2, there exists, for each $2\leq n<\infty$, such a pair $K\leq Q$ where $Q$ is finitely presented. Therefore one can use the Rips construction [Rip82] or any of its variants (see e.g. [Are23] and the references therein), to produce examples of finitely generated pairs $H\leq G$ where $H\operatorname{\setminus}G$ is an $n$-ended quasi-tree, $H$ is not a virtual fiber and $G$ is, in some precise way, “well-behaved”. For example, Arenas’ construction [Are23] yields a group $G$ which is torsion-free hyperbolic (in particular, of type $F$) and virtually compact special.